# Enhancing Temporal Sensitivity and Reasoning for Time-Sensitive Question Answering

### v1

@@ Abstract
Wanqi Yang1∗,  Yanda Li1thanks: Equal contributions,  Meng Fang2,  Ling Chen1thanks: Corresponding Author 1 University of Technology Sydney  2 University of Liverpool 1wanqi.yang-1@student.uts.edu.au, 1liyanda95@gmail.com 2Meng.Fang@liverpool.ac.uk, 1ling.chen@uts.edu.au
@@ 1 Introduction
Time-Sensitive Question Answering (TSQA) involves parsing and responding to questions that depend on specific time points or periods. For instance, Obama’s roles in 2006 and 2016 were distinctly different. Unlike conventional Question Answering (QA) tasks rajpurkar2016squad; joshi2017triviaqa; dunn2017searchqa; ye2023multi; zhao2023verify, TSQA requires language models to discern and understand temporal information within questions. Moreover, it necessitates the capacity to locate relevant facts using temporal information, leveraging time-related knowledge to provide accurate and relevant answers. In the domain of TSQA, models are required to predict answers through generation or extraction methods, given a set of questions and contexts embedded with temporal information. These questions are characterized by explicit or implicit temporal expressions, while the context invariably contains multiple facts evolving over time. Recently, several TSQA datasets have been introduced, including notable work chen2021dataset that provides 40k time-sensitive questions with corresponding free-text contexts derived unprocessed from WikiData vrandevcic2014wikidata. Building upon this foundation,  tan2023towards not only furnish free-text contexts from WikiData without any modification but also offer a version of the dataset where the context has been restructured. These contributions have facilitated evaluations of existing large language models (LLMs) such as FiD izacard2020leveraging, BigBird zaheer2020big, and T5 raffel2020exploring on these datasets, underscoring the challenge TSQA poses, as the performance of LLMs remains significantly inferior to human levels. However, an effective methodology to significantly improve LLMs performance in TSQA is still lacking.
It is challenging to improve model’s sensitivity to temporal information and the capacity for temporal reasoning. Firstly, LLMs significantly lack attention to and comprehension of temporal information in both questions and their contexts ning2020torque; shang2021open. It is essential for models to prioritize temporal cues in questions to accuately locate corresponding times and facts within contexts, especially for lengthy and complex free-text contexts, which encompass multiple irrelevant information and various chronological connections. Focusing on temporal information allows a model to circumvent the distraction of irrelevant data and to swiftly locate relevant facts through the chronological connections among different paragraphs. Secondly, LLMs han2020econet; dhingra2022time exhibit notably weak temporal reasoning abilities. This specifically refers to the models’ susceptibility to interference from answers that pertain to the same entity or relationship but span different time periods within the context, as well as to disturbances from other events occurring within the same time period.
@@ 3.1 Temporal Information-Aware Embedding
1) Firstly, we construct a question temporal matrix $A_{q}$ and a context temporal matrix $A_{c}$, both initially filled with zeros.
@@
In our time-aware module, we use SpaCy to locate temporal expressions from question $a_{q_{t}}$ and context $a_{c_{t}}$, marking these detected positions in $A_{q}$ and $A_{c}$ with ones. SpaCy’s entity recognition module to identify a variety of time entities, including but not limited to: 1) Complete dates (January 1, 2020; 2020/01/01) 2) Month and year (January 2020) 3) Season and year (Spring 2021) 4) Month and day (1 January) 5) Year (2020) 6) Weekdays 7) Relative times (next week, last year) 8) Holidays (Christmas, New Year’s Day), etc.
@@
2) Employing a sliding window $W$ with window size $L$ centered on the temporal information, we use $f(W)$ function to mark the positions within the window region as ones to further highlight regions of temporal relevance.
@@ 3.2 Granular Contrastive Reinforcement Learning
we propose a novel contrastive reinforcement learning framework, building on previous methods like the EM algorithm tan2023towards, which occasionally fails to recognize semantically equivalent responses due to its rigid scoring mechanics. For instance, when the model’s predicted answer is “the capital of France” and the ground truth is “Paris”, the EM algorithm assigns a score of 0. Nonetheless, it is widely acknowledged that “the capital of France” and “Paris” should be considered equivalent answers.
@@
This approach not only refines the model’s ability to make temporally precise judgments but also enhances its overall learning efficiency by dynamically adjusting to the complexities of temporal reasoning. Thereby significantly boosting model’s learning efficiency and effectiveness. The model is optimized using Proximal Policy Optimization schulman2017proximal, whose details are shown in Appendix A.
@@ 4.1 Experimental Setup
  • Time-Event QA (L2) is derived from the TempReason  tan2023towards dataset. The distinguishing features of L2 include: 1) The presence of a specific time point in each question, denoted by “in YEAR”; 2) The time point mentioned in the question may not find a direct match within the context; 3) Context is provided from Wikidata (OBQA); 4) A structured context is offered, where the content is strictly organized according to the ($s$, $r$, $o$, [$t_{s}$, $t_{e}$]) template (ReasonQA).
@@
  • TimeQA Easy is based on the TimeQA dataset  chen2021dataset. Features of the Easy Version include: 1) Each question contains a time expression either in the format “in YEAR” or “from YEAR to YEAR”; 2) The time mentioned can be directly matched within the context; 3) Context from Wikidata is provided (OBQA).
@@
In the first stage of training , we train T5-base model for 6 epochs with a batch size of 8 on a NVIDIA Tesla V100 GPUs. The model is optimized using AdamW loshchilov2017decoupled with a learning rate of $5e^{-6}$. In the second stage of training, we fine-tune the model with our framework for 10 epochs with a batch size of 16. The model is optimized using Proximal Policy Optimization schulman2017proximal. In our configuration of the Temporal Information-Aware Embedding, we employ SpaCy, an open-source Natural Language Processing (NLP) library, as time-aware module to locate temporal information within the provided question and context. The window size of the sliding window is set to 10. Furthermore, in Granular Contrastive Reinforcement Learning, the ratio between remote negative answers and proximal negative answers is established at 1:1.
@@
  • FLAN-T5-Large wei2021finetuned represents an advanced instantiation of NLP models based on the extensive T5 framework. We have deployed FLAN-T5-Large for the purpose of evaluating its performance across a variety of TSQA datasets.
  • ChatGPT ouyang2022training, developed by OpenAI, is a sophisticated NLP model predicated on the GPT (Generative Pre-trained Transformer) architecture. Utilizing the official API provided for gpt-3.5-turbo, we have conducted evaluation on various TSQA datasets.
  • T5-SFT raffel2020exploring, an acronym for “Text-to-Text Transfer Transformer”, is a NLP model conceived by the Google Research team. We subjected the TSQA dataset to supervised fine-tuning on the T5-base model to conduct further evaluation.
@@ 4.2 Experimental Results
Table 5: Comparison with TempT5 tan2023towards.
@@ 4.3 Ablation Study
Furthermore, we conducted comparative experiments with another framework named TempT5 tan2023towards specifically designed for the TempReason dataset, as shown in Table 5. Our framework demonstrated a markedly superior performance.
@@ 8 Acknowledgements
We would like to express our gratitude for the support provided by the Australian Research Council Discovery Project, grant number ARC DP240101349, which partially supported this research.

### v2

@@ Abstract
Wanqi Yang1∗,  Yanda Li1thanks: Equal contributions,  Meng Fang2,  Ling Chen1 1 University of Technology Sydney  2 University of Liverpool 1wanqi.yang-1@student.uts.edu.au, 1Yanda.Li@student.uts.edu.au 2Meng.Fang@liverpool.ac.uk, 1ling.chen@uts.edu.au
@@ 1 Introduction
Time-Sensitive Question Answering (TSQA) involves parsing and responding to questions that depend on specific time points or periods. For instance, Obama’s roles in 2006 and 2016 were distinctly different. Unlike conventional Question Answering (QA) tasks Rajpurkar et al. (2016); Joshi et al. (2017); Dunn et al. (2017); Ye et al. (2023); Zhao et al. (2023), TSQA requires language models to discern and understand temporal information within questions. Moreover, it necessitates the capacity to locate relevant facts using temporal information, leveraging time-related knowledge to provide accurate and relevant answers. In the domain of TSQA, models are required to predict answers through generation or extraction methods, given a set of questions and contexts embedded with temporal information. These questions are characterized by explicit or implicit temporal expressions, while the context invariably contains multiple facts evolving over time. Recently, several TSQA datasets have been introduced, including notable work Chen et al. (2021) that provides 40k time-sensitive questions with corresponding free-text contexts derived unprocessed from WikiData Vrandečić and Krötzsch (2014). Building upon this foundation,  Tan et al. (2023) not only furnish free-text contexts from WikiData without any modification but also offer a version of the dataset where the context has been restructured. These contributions have facilitated evaluations of existing large language models (LLMs) such as FiD Izacard and Grave (2020), BigBird Zaheer et al. (2020), and T5 Raffel et al. (2020) on these datasets, underscoring the challenge TSQA poses, as the performance of LLMs remains significantly inferior to human levels. However, an effective methodology to significantly improve LLMs performance in TSQA is still lacking.
It is challenging to improve model’s sensitivity to temporal information and the capacity for temporal reasoning. Firstly, LLMs significantly lack attention to and comprehension of temporal information in both questions and their contexts Ning et al. (2020); Shang et al. (2021). It is essential for models to prioritize temporal cues in questions to accuately locate corresponding times and facts within contexts, especially for lengthy and complex free-text contexts, which encompass multiple irrelevant information and various chronological connections. Focusing on temporal information allows a model to circumvent the distraction of irrelevant data and to swiftly locate relevant facts through the chronological connections among different paragraphs. Secondly, LLMs Han et al. (2020); Dhingra et al. (2022) exhibit notably weak temporal reasoning abilities. This specifically refers to the models’ susceptibility to interference from answers that pertain to the same entity or relationship but span different time periods within the context, as well as to disturbances from other events occurring within the same time period.
@@ 3.1 Temporal Information-Aware Embedding
1) Firstly, we construct a question temporal matrix $A_{q}$ and a context temporal matrix $A_{c}$, both initially filled with zeros as:
@@
In our time-aware module, we use SpaCy to locate temporal expressions from question $a_{q_{t}}$ and context $a_{c_{t}}$, marking these detected positions in $A_{q}$ and $A_{c}$ with ones:
@@
2) Employing a sliding window $W$ with window size $L$ centered on the temporal information, we use $f(W)$ function to mark the positions within the window region as ones to further highlight regions of temporal relevance. The definition of $W$ for an element $a_{i}$ is given by:
@@ 3.2 Granular Contrastive Reinforcement Learning
we propose a novel contrastive reinforcement learning framework, building on previous methods like the EM algorithm Tan et al. (2023), which occasionally fails to recognize semantically equivalent responses due to its rigid scoring mechanics. For instance, when the model’s predicted answer is “the capital of France” and the ground truth is “Paris”, the EM algorithm assigns a score of 0. Nonetheless, it is widely acknowledged that “the capital of France” and “Paris” should be considered equivalent answers.
@@
This approach not only refines the model’s ability to make temporally precise judgments but also enhances its overall learning efficiency by dynamically adjusting to the complexities of temporal reasoning. Thereby significantly boosting model’s learning efficiency and effectiveness. The model is optimized using Proximal Policy Optimization Schulman et al. (2017), whose details are shown in Appendix A.
@@ 4.1 Experimental Setup
  • Time-Event QA (L2) is derived from the TempReason  Tan et al. (2023) dataset. The distinguishing features of L2 include: 1) The presence of a specific time point in each question, denoted by “in YEAR”; 2) The time point mentioned in the question may not find a direct match within the context; 3) Context is provided from Wikidata (OBQA); 4) A structured context is offered, where the content is strictly organized according to the ($s$, $r$, $o$, [$t_{s}$, $t_{e}$]) template (ReasonQA).
@@
  • TimeQA Easy is based on the TimeQA dataset  Chen et al. (2021). Features of the Easy Version include: 1) Each question contains a time expression either in the format “in YEAR” or “from YEAR to YEAR”; 2) The time mentioned can be directly matched within the context; 3) Context from Wikidata is provided (OBQA).
@@
In the first stage of training , we train T5-base model for 6 epochs with a batch size of 8 on a NVIDIA Tesla V100 GPUs. The model is optimized using AdamW Loshchilov and Hutter (2017) with a learning rate of $5e^{-6}$. In the second stage of training, we fine-tune the model with our framework for 10 epochs with a batch size of 16. The model is optimized using Proximal Policy Optimization Schulman et al. (2017). In our configuration of the Temporal Information-Aware Embedding, we employ SpaCy, an open-source Natural Language Processing (NLP) library, as time-aware module to locate temporal information within the provided question and context. The window size of the sliding window is set to 10. Furthermore, in Granular Contrastive Reinforcement Learning, the ratio between remote negative answers and proximal negative answers is established at 1:1.
@@
  • FLAN-T5-Large Wei et al. (2021) represents an advanced instantiation of NLP models based on the extensive T5 framework. We have deployed FLAN-T5-Large for the purpose of evaluating its performance across a variety of TSQA datasets.
  • ChatGPT Ouyang et al. (2022), developed by OpenAI, is a sophisticated NLP model predicated on the GPT (Generative Pre-trained Transformer) architecture. Utilizing the official API provided for gpt-3.5-turbo, we have conducted evaluation on various TSQA datasets.
  • T5-SFT Raffel et al. (2020), an acronym for “Text-to-Text Transfer Transformer”, is a NLP model conceived by the Google Research team. We subjected the TSQA dataset to supervised fine-tuning on the T5-base model to conduct further evaluation.
@@ 4.2 Experimental Results
Table 5: Comparison with TempT5 Tan et al. (2023).
@@ 4.3 Ablation Study
Furthermore, we conducted comparative experiments with another framework named TempT5 Tan et al. (2023) specifically designed for the TempReason dataset, as shown in Table 5. Our framework demonstrated a markedly superior performance.
5 Related Work
Time Sensitive Question Answering
Recent developments in Open Book Question Answering (OBQA) have brought about significant advancements in addressing time-sensitive questions. These advancements are particularly evident in contemporary Time Sensitive Question Answering (TSQA) datasets  Zhang and Choi (2021); Chen et al. (2021); Tan et al. (2023); Yang et al. (2024).
SituatedQA Zhang and Choi (2021) stands out as a notable contribution to the field. This dataset focuses on open-domain time-sensitive QA, featuring realistic questions reannotated from the NQ dataset  Kwiatkowski et al. (2019) to incorporate context dependence and diverse answers across temporal or geographical contexts.
Similarly, TimeQA Chen et al. (2021) presents a dataset comprising 20K questions with an emphasis on time sensitivity. Notably, TimeQA’s hard version challenges models to reason over implicit temporal mentions within passages, a feature less emphasized in SituatedQA.
Tan’s work Tan et al. (2023) in this domain is exemplified by the development of the TempReason benchmark for TSQA. This benchmark offers a comprehensive assessment covering various types of temporal understanding, including time-time (L1) relations, time-event (L2) relations, and event-event (L3) relations.
These TSQA datasets adopt the OBQA setting, which leverages external context, such as natural language text, to assist Language Models (LMs) in answering questions. This contrasts with Closed Book Question Answering (CBQA), where only the question is provided to the LM without access to external text  Févry et al. (2020); Roberts et al. (2020); Dhingra et al. (2022); Liska et al. (2022); Wei et al. (2023).
Overall, TSQA tasks highlight the complexity and realism inherent in time-sensitive QA, emphasizing the necessity for LMs to comprehend temporal relationships and effectively ground information temporally to provide accurate responses.
Large Language Models on TSQA
Some studies have shown that large language models (LLMs) such as those developed by Devlin et al. (2018), Raffel et al. (2020), and Liu et al. (2019) exhibit strong performance in general question answering tasks Rajpurkar et al. (2016); Kwiatkowski et al. (2019). However, with the emergence of Time-Sensitive Question Answering (TSQA) tasks and related datasets, LMs have demonstrated less satisfactory performance. For instance,  Chen et al. (2021) demonstrated that large language models like FiD Izacard and Grave (2020) and BigBird Zaheer et al. (2020), functioning as generative and extractive QA models, achieve only 60% accuracy on TSQA. Furthermore, experimental results from  Tan et al. (2023) showed that FLANT5-Large Wei et al. (2021) and ChatGPT Ouyang et al. (2022) achieve performance of only around 10%.
To address the challenges of TSQA,  Jia et al. (2018b) proposed TEQUILA, a method specifically tailored for temporal knowledge graph-based QA. TEQUILA utilizes constraint reasoning on temporal intervals to compute answers to question.  Faghihi and Kordjamshidi (2021) introduced the Time-stamped Language Model as a novel approach to comprehend the progression of events. Their work extends the understanding of event flows across time. Additionally, Inspired by work on relational KGQA Huang et al. (2019); Saxena et al. (2020),  Shang et al. (2022) proposed a method of improving time sensitivity for question answering over temporal knowledge graphs Jia et al. (2018a, 2021); Saxena et al. (2021), which incorporates contrastive learning, time-sensitive temporal KG embedding, and KG-based pruning techniques.  Mavromatis et al. (2022) introduced a novel framework by proposing a joint model which integrates temporal knowledge graph embeddings with pre-trained language models. These methods have shown excellent performance on Temporal Knowledge Graphs data. However, as far as we know, there is a lack of effective methods for handling free-text TSQA datasets.
@@ 9 Acknowledgements
This project is partially supported by ARC DP240101349.